# Predictability-Aware Compression and Decompression Framework for Multichannel Time Series Data

### v1

@@ 2.1 Compression process
To demonstrate that the encoding process uncovers the predictability of time series data, a justification is provided based on the Pearson correlation coefficient. The corresponding result is shown below (The detailed derivation is provided in Section A.1.):
$R$ $=\frac{\sum_{t}^{t+\tau}(y_{t}^{i}-\bar{y}^{1})(y_{t+\tau}^{i}-\bar{y}^{2})}{\sqrt{\sum_{t}^{t+\tau}(y_{t}^{i}-\bar{y}^{1})^{2}}\sqrt{\sum_{t}^{t+\tau}(y_{t+\tau}^{i}-\bar{y}^{2})^{2}}}$ (2)
$=\frac{\overline{\gamma}[S_{1}^{2}+...+S_{\tau}^{2}]}{\left|\overline{\gamma}\right|[S_{1}^{2}+...+S_{\tau}^{2}]}$
$=1$
where $\overline{\gamma}>0$, and $S_{i}$ is computed in the following formulation,
$S_{t}=[y_{t}^{i}-y_{\text{mean}}^{\tau}]^{2}$ (3)
Multichannel combination
Following the principle of Least Common Multiple Period (LCMP) in signal processing, if $T_{1}$ and $T_{2}$ are the fundamental periods of two periodic signals, then their least common multiple $\mathrm{lcm}(T_{1},T_{2})$ is typically the fundamental period of the combined signal $z(t)$. To maintain inter-channel period consistency and combined representation predictability, all seasonal keys are constructed with the same period $\tau=\mathrm{lcm}(\tau_{1},\tau_{2},\dots,\tau_{C})$. Therefore, for multichannel compression, each channel is first individually encoded using its corresponding periodic key, and the encoded representations are then combined. The overall compression process is shown below:
$Y$ $=\sum_{i=1}^{C}Y^{i}$ (4)
$=\sum_{i=1}^{C}\text{Encode}(K,X^{i})$
where $C$ is the number of channels.
Compression optimization
Although the complexity of the compression process does not dominate the overall time complexity in the analysis (shown in 2.3), a sparse compression strategy is introduced to further reduce the computational burden on edge devices. Concretely, the sparse compression process is formulated as follows:
$Y^{C}$ $=\sum_{i=0}^{C}[Y_{0:\tau}^{i},Y_{\tau:2\tau}^{i},...,Y_{(\lfloor\frac{N}{\tau}\rfloor-1)\tau:\lfloor\frac{N}{\tau}\rfloor\tau}^{i}]$ (5)
$=\sum_{i=0}^{C}[\text{Encode}(X_{0:\tau}^{i},K),\text{Encode}(X_{\tau:2\tau}^{i},K),...,\text{Encode}(X_{(\lfloor\frac{L}{\tau}\rfloor-1)\tau:\lfloor\frac{L}{\tau}\rfloor\tau}^{i},K)]$
[…]
$\mathcal{O}(N)_{\text{compression}}=\frac{CN^{2}}{\tau}$ (6)
When $\tau=\frac{N}{\log N}$, the time complexity reduces to $\mathcal{O}(CN\log N)$.
2.2 Decompression process
The decompression process is designed to mitigate the additional reconstruction errors caused by the oversimplified assumption in HRR, which assumes that the input data follow a normal distribution—an assumption that is often inappropriate for time series data. In this section, we first present the decoding process based on circular correlation, including the analysis of reconstruction errors. We then introduce our proposed method to alleviate these errors.
Decoding process
Motivated by HRR, the decoding process using circular correlation process and periodic keys is defined as follows:
$\tilde{X}^{i}$ $=\text{Decode}(K,Y)$ (7)
$=[\sum_{z=0}^{n-1}y_{z}k_{z-0},\sum_{z=0}^{n-1}y_{z}k_{z-1},...,\sum_{z=0}^{n-1}y_{z}k_{z-{h}}]$
$=[(\sum k^{2})(x_{0}^{0}+x_{0}^{1}+...+x_{0}^{l-1})+\eta,(\sum k^{2})(x_{1}^{0}+x_{1}^{1}+...+x_{1}^{C-1})+\eta$
$,...,(\sum k^{2})(x_{h}^{0}+x_{h}^{1}+...+x_{h}^{C-1})+\eta]$
where $\eta$ denotes the accumulated intra-channel errors resulting from different timestamps within the same channel. HRR assumes that both the data $X$ and the key $K$ follow a normal distribution, which allows intra-channel interference to be reduced. However, this assumption does not hold for time series data and periodic keys, thereby motivating the design of a new decompression process. In addition to intra-channel errors, recovering the $i$-th channel result $\tilde{X}^{i}$ also suffers from inter-channel interference introduced by other channels at the same timestamp, further degrading reconstruction quality.
Intra-channel errors reduction
To reduce the effect of intra-channel errors, we adopt an orthogonal circular matrix for keys $K$. The characteristic of a circular matrix [5] is that each column is a cyclic shift of the previous column, meaning it is entirely determined by its first row. An orthogonal circular matrix is a special type of circular matrix that satisfies the orthogonality condition $K^{T}K=I$, where $I$ is the identity matrix. (Obtained from two times FFT [24]).
Based on the properties of the orthogonal circular keys Matrix, different shifts (or permutations) of the key vectors are mutually orthogonal. During the decoding process, when recovering a specific timestamp, the multiplication of the corresponding key vector with those from other timestamps results in zero due to the orthogonality. Consequently, only the current timestamp’s contribution remains when decoded using the key matrix (see Appendix A.2 for details). The updated reconstruction based on the orthogonal circular key matrix is given by:
$\tilde{X}^{i}$ $=(\sum k^{2})\cdot\begin{bmatrix}x_{0}^{0}+x_{0}^{1}+...+x_{0}^{C-1}\\ x_{1}^{0}+x_{1}^{1}+...+x_{1}^{C-1}\\ ...\\ x_{n-1}^{0}+x_{n-1}^{1}+...+x_{n-1}^{C-1}\end{bmatrix}^{T}$ (8)
Inter-channel errors reduction
To eliminate inter-channel interference and restore channel-wise fidelity, we introduce a residual correction module. The main idea is that utilize the residual process to adaptively remove the other channel’s effect on the same timestamps. The process is shown below,
@@ 2.2 Decompression process
$\tilde{X}=\text{Dense}(\frac{\text{Copy}(\text{Decode}(K,Y))}{\sum k^{2}}+\mathrm{Residual}(\frac{\text{Copy}(\text{Decode}(K,Y))}{\sum k^{2}}))\$ (9)
where the $\mathrm{Residual}(\cdot)$ denotes a lightweight residual block consisting of two convolution layers with one ReLU activation, and $\text{Copy}()$ denotes either duplicating the decoded result to match the number of channels or using a learnable transformation (e.g., a dense layer) to project it to the desired channel dimension. This correction allows the model to disentangle individual channels from the shared reconstruction, thereby enhancing decompression precision. The associated time complexity is given in Equation 11.
2.3 Time complexity analysis
$\mathcal{O}(N)_{\text{comp}}$ $=\underbrace{LTN^{2}}_{\text{Prediction}}+\underbrace{\frac{CN^{2}}{\tau}+2N\log N}_{\text{Compression}}+\underbrace{N+3NC+NC^{2}+CN^{2}}_{\text{Decompression}}$ (10)
[…]
where $T$ denotes the number of training epochs, $N$ is the number of timestamps, $C$ $(C<N)$ is the number of channels, $L$ is the depth of the predictor, and $\tau$ is the least common multiple (LCM) period across different channels. The time complexity of the residual block in the decompression process is given by:
$\mathcal{O}(N)_{\text{decompression}}$ $=\mathcal{O}(\text{Divide})+\mathcal{O}(\text{Conv})+\mathcal{O}(\text{Relu})+\mathcal{O}(\text{Conv})+\mathcal{O}(\text{Residual})+\mathcal{O}(\text{Dense})$ (11)
$=N+NC+NC+NC^{2}+NC+CN^{2}$
$=N+3NC+NC^{2}+CN^{2}$
In contrast, a naive multichannel predictor without compression incurs the following computational complexity:
$\mathcal{O}(N)_{\text{naive}}=LTCN^{2}.$ (12)
To determine the minimum number of channels $C$ for which the proposed method achieves lower computational cost, the two complexities are compared:
$\mathcal{O}(N)_{\text{naive}}$ $>\mathcal{O}(N)_{\text{comp}}$ (13)
$LTCN^{2}$ $>LTN^{2}+CN^{2}+\frac{CN^{2}}{\tau}$
$\Rightarrow\quad C$ $>\frac{LT}{LT-1-\frac{1}{\tau}}\quad\text{(assuming }LT\gg 1\text{ and }\frac{1}{\tau}\to 0\text{)}$
Under the assumptions $LT\gg 1$ and $\frac{1}{\tau}\to 0$, the lower bound reduces to $C>1$, indicating that the proposed framework offers a computational advantage when the number of channels exceeds one. Furthermore, since the overhead introduced by compression and decompression is subdominant in the asymptotic regime, it does not alter the scaling advantage of the proposed method.
2.4 Training strategy
Normalization
Following Dish-TS [7], a standard normalization is applied after compression and reversed before decompression. Specifically, each compressed segment is normalized by subtracting its mean and dividing by its standard deviation to stabilize outputs across different periods $\tau$. During inference, the predicted outputs are denormalized accordingly to restore the original scale.
Training and testing process
The total training objective is composed of three components:
$\mathcal{L}$ $=\mathcal{L}_{\text{Prediction}}+\mathcal{L}_{\text{Regulation}}+\mathcal{L}_{\text{Normal}}$ (14)
$=(X_{\text{Label}}-X_{\text{Predict}})^{2}+\alpha(\sum_{1}^{C}|X_{\text{Residual}}|-\sum_{1}^{C}X_{\text{Predict}})^{2}+\beta(\varphi_{\text{Y}}-\varphi_{\hat{Y}})^{2}$
Here, $\mathcal{L}_{\text{Prediction}}$ denotes the prediction error, $\mathcal{L}_{\text{Regulation}}$ denotes the regulation term for the decompression process, $\mathcal{L}_{\text{Normal}}$ denotes the mean consistency between the predicted andd true signals.
To enhance the numerical stability during the training of the compression–prediction–decompression framework, we adopt a scale-invariant gradient clipping strategy [14]. Specifically, given a parameter $w$ and its corresponding gradient $g$, we modify the gradient as follows:
$g\leftarrow g\times\min(1,\frac{\alpha\|w\|}{\|g\|}),$ (15)
where $\alpha>0$ is a hyperparameter controlling the clipping threshold. This approach adaptively scales the gradient based on the magnitude of the parameter, ensuring that the update is proportional to the parameter norm. Such scale-invariant clipping prevents exploding gradients without introducing excessive regularization, and is particularly effective in architectures with large scale variations across different layers (see Appendix A.5 for details).
3 Experiments
3.1 Experimental settings
Datasets
To demonstrate the effectiveness of our predictability-aware compression and decompression framework (PCDF) in multichannel time series prediction, we evaluate it on a diverse set of application datasets collected from edge device sensors. These datasets include traffic data (NYC Taxi [19], DC Bike [1]), energy data (solar energy generation [18], electricity consumption [28]), weather data [3], and manufacturing data (gas sensor array drift [30]).
Baselines and setup
We apply our compression–decompression framework to multiple existing predictors (MLP, Transformer, Linear, and Convolution, as detailed in Appendix A.7). Throughout the paper, we refer to each such combination as an integration (e.g., PCDF-MLP, PCDF-Trans, PCDF-Linear, PCDF-Conv). We compare these integrations with state-of-the-art models, including PatchTST [21], Autoformer [31], HDMixer [10], DLinear [32], MSGNet [2], and TSMixer [6], to evaluate changes in performance and runtime trade-offs. All models follow the same experimental setup, with the number of prediction channels set to $C\in\{5,10,20,30,40\}$. All experiments are conducted on two NVIDIA RTX 4090 GPUs, each equipped with 24 GB of memory.
3.2 Prediction MSE and runtime with Comprehensive analysis
@@
Table 1: Cobb–Douglas production index (lower values indicate better overall performance)
@@
Following the idea of the Cobb–Douglas production function [4], which uses a multiplicative combination to model the joint contribution of multiple factors, we combine MSE and computational time in a similar fashion to capture the trade-off and enable a comprehensive analysis of prediction accuracy and computational efficiency.
Based on the MSE and runtime results for all models (see details in Appendix A.8 and Appendix A.9), the multiplicative combination of MSE and runtime—referred to as the Cobb–Douglas production index—is presented in Table 1 and formally defined in Equation 16. The compression–decompression integrations achieve lower combined index values compared to the representative baseline models across different datasets in most cases. These results suggest that the compression–decompression framework has the potential to achieve better overall performance with reduced prediction time and lower prediction errors.
[…]
From the perspective of different datasets, those with clear periodic patterns, such as Electricity, tend to exhibit lower combined index values. This indicates that such datasets either require less computational time or result in lower prediction errors after processing, leading to better overall performance. Conversely, datasets with less pronounced periodic patterns tend to have higher combined index values, reflecting increased prediction errors or longer computational times.
3.3 Variance analysis of MSE, runtime, and composite index across datasets
Figure 3 shows the variance of MSE and runtime. The integrations exhibit lower variance across different datasets compared to their corresponding base models, indicating that our framework provides better stability for the underlying predictors.
@@ 3.3 Variance analysis of MSE, runtime, and composite index across datasets
Figure 3: Boxplot of MSE and runtime across channel configurations and datasets
In terms of value distribution, the integrations generally achieve lower values than their base models, except for the PCDP-MLP and its corresponding base model TSMixer in the MSE metric. This suggests that our framework can effectively reduce runtime, decrease MSE, and improve overall performance compared to other methods built upon the same predictors—despite the exception observed with TSMixer.
3.4 Ablation results
Table 2: Compression-decompression framework ablation study via MSE
@@ 3.4 Ablation results
Compression-decompression 0.1799515 0.215305 0.10726 0.2172095 0.238878
Randomkey-decompression 0.19826 0.259521 14.803542 18.559643 0.253261
Compression-decoder 0.213246 1257.967764 0.2435455 9.832429 0.5709225
Encoder-decompression 0.323226 5.4924415 143.518735 1.998019 12.7464045
Encoder-decoder 0.1545935 1.5364435 0.173675 1.6343485 0.302809
In the ablation study, a classic encoder–decoder framework is used to replace the compression and decompression components in order to justify the efficiency of the proposed framework. In addition, a random key is used in place of the periodic key to evaluate the effectiveness of the compression component guided by periodic information. Therefore, the ablation study includes four variants: (1) random-key compression–decompression, (2) compression–decoder, (3) encoder–decompression, and (4) encoder–decoder.
In the random-key variant, the results vary significantly across datasets, suggesting that the random key fails to uncover the underlying predictive patterns in some datasets. This highlights the importance of using the periodic key within the compression–decompression framework.
For the compression–decoder and encoder–decompression variants, since compression and decompression are co-designed and interdependent, using only one of the components leads to unstable results. This indicates that optimal performance relies on the synergy between both compression and decompression modules.

### v2

@@ 5.2. Decompression
$=\left[\sum_{z=0}^{n-1}\hat{y}_{z}k_{z},\ \sum_{z=0}^{n-1}\hat{y}_{z}k_{z-1},\ \dots,\ \sum_{z=0}^{n-1}\hat{y}_{z}k_{z-h}\right],$
where $\hat{Y}=[\hat{y}_{0},\hat{y}_{1},\dots,\hat{y}_{h-1}]$ is the predicted representation from the forecasting model, $h$ denotes the length of the predicted future sequence to be reconstructed, and $k_{z}$ is the $z$-th element of the decoding kernel $K$. Under HRR, this yields:
(9) $\tilde{X}^{i}[t]\approx(\sum k^{2})\cdot\sum_{c=1}^{C}x_{t}^{c}+\eta_{t},$
where $\eta_{t}$ denotes accumulated intra-channel interference due to overlapping shifts in the circular convolution.
Reducing intra-channel errors via orthogonal keys.
To reduce intra-channel interference, we use orthogonal circular key matrices $K$, where each key shift is orthogonal to others: $K^{\top}K=I$. Such matrices can be efficiently constructed via two FFTs (Wang and others, 2018). In this setting, different time shifts become mutually orthogonal, nullifying off-timestamp interference in decoding:
(10) $\tilde{X}^{i}[t]=(\sum k^{2})\cdot\sum_{c=1}^{C}x_{t}^{c}$
Inter-channel error correction via residuals.
While orthogonality mitigates intra-channel interference, inter-channel errors still remain. To address this, we introduce a residual correction block. The final decompression is expressed as:
@@
(11) $\hat{X}=\text{Dense}\left(\frac{\text{Copy}(\tilde{X})}{\sum k^{2}}+\mathrm{Residual}\left(\frac{\text{Copy}(\tilde{X})}{\sum k^{2}}\right)\right)$
where $\tilde{X}=\text{Decode}(K,Y)$ with orthogonal circular key matrices, $\text{Copy}(\cdot)$ duplicates or projects the decoded result into the full channel dimension, $\mathrm{Residual}(\cdot)$ denotes a lightweight correction block (2 conv layers + ReLU), $\text{Dense}(\cdot)$ is fully connected layer. This decomposition corrects for shared-channel entanglement and significantly enhances decompression fidelity.
5.3. Time Complexity Superiority
This part provides a time complexity analysis for compression and decompression framework.
Theorem 3: Computational Superiority Condition.
Let $\mathcal{O}(L)_{\text{comp}}$ denote the total computational complexity of our compression framework followed by a single-channel predictor, and let $\mathcal{O}(L)_{\text{naive}}$ denote that of a naive multichannel predictor, which shares the same structure as the single-channel predictor except that it operates on all channels directly. Here, $E$ denotes the number of training epochs and $D$ denotes the depth of predictor. Then:
(12) $\mathcal{O}(L)_{\text{comp}}<\mathcal{O}(L)_{\text{naive}}\\ \Rightarrow C>\frac{DE}{DE-1-\frac{1}{\tau}}.$
5.3.1. Corollary 2: Simplified Superiority Condition.
Under the asymptotic regime $DE\gg 1$ and $\frac{1}{\tau}\to 0$, the condition reduces to $C>1$. The proof of Proposition 4 and Corollary is provided in Appendix A.3.
Remark 5
Theorem 3 and Corollary 2 together indicate that the proposed compression-decompression framework offers a computational advantage over the naive multichannel predictor when the number of input channels $C$ exceeds one. Furthermore, the overhead introduced by compression and decompression becomes subdominant. As a result, the scaling advantage of the proposed framework remains intact in practical high-dimensional time series settings.
6. Experiments
In this section, we first introduce the experimental settings, including datasets, baselines, and evaluation metrics. We then analyze the results of the comparative experiments from different metrics aspect.
@@ 6. Experiments
Table 2. Cobb–Douglas Production Index (CDPI) results. Lower values indicate better overall performance. Underlined values denote the best performance in each setting. The number of channels corresponds to the compression ratio, as our framework compresses all channels into one.
@@
6.1. Experimental Settings
Datasets.
We evaluate our design on datasets from a diverse set of edge-based sensing applications, including (1) traffic: NYC taxi (New York City Taxi and Limousine Commission, 2025) and DC_bike (Capital Bikeshare, 2025); (2) energy: solar energy generation (National Renewable Energy Laboratory, 2025) and electricity consumption (Trindade, 2015); (3) manufacturing: gas sensor array drift (Vergara and others, 2012); and (4) weather (Cejpek and contributors, 2025).
Baselines and setup.
We implement our framework with multiple existing predictors (MLP, Transformer, Linear, and Convolution, as detailed in the Appendix). Each such combination is referred to as an integration (e.g., PCDF-MLP, PCDF-Trans, PCDF-Linear, PCDF-Conv). The architectural details of each single-channel predictor are described as follows:
MLP integration: Built upon the TSMixer-based MLP, this variant incorporates an additional flatten layer, one linear layer, and a ReLU activation, appended to the end of the original structure.
Transformer integration: Based on the PatchTST encoder, this variant extends the original architecture by adding a flatten layer, two linear layers, and a ReLU activation.
Linear integration: A linear predictor derived from HDMixer, consisting of multiple linear layers, with a flatten operation inserted before a final fully connected layer.
Convolution integration: A convolutional module comprising two convolutional layers, followed by a flatten operation, two fully connected layers, and a ReLU activation function.
We compare these integrations with state-of-the-art models, including PatchTST (Nie and others, 2023), Autoformer (Wu and others, 2021), HDMixer (Huang and others, 2024), DLinear (Zeng and others, 2023), MSGNet (Cai and others, 2024), and TSMixer (Ekambaram and others, 2023), to evaluate performance and runtime trade-offs. Since, to the best of our knowledge, there is no existing multichannel compression framework specifically designed for time series prediction (as discussed in the Introduction and Related Work), no direct baselines are available for this setting. All models follow the same experimental setup, with the number of prediction channels set to $C\in\{5,10,20,30,40\}$. All experiments are conducted on two NVIDIA RTX 4090 GPUs, each equipped with 24 GB of memory.
@@ 6.1. Experimental Settings
Figure 4. Boxplot of MSE and runtime across channel configurations and datasets
Metric.
Evaluation metrics include Mean Squared Error (MSE), runtime, and the Cobb–Douglas Production Index (CDPI). Runtime is obtained from the testing process, not from the training process; therefore, the number of training epochs does not affect the experimental results.
CDPI follows the idea of the Cobb–Douglas production function (Cobb and others, 1928), which uses a multiplicative combination to model the joint contribution of multiple factors. We combine MSE and runtime in a similar fashion to capture the trade-off and enable a comprehensive analysis of prediction accuracy and computational efficiency.
(13) $\text{CDPI}=\text{MSE}\cdot\text{Runtime}.$
In our experiments, we do not apply normalization when computing CDPI, in order to preserve the raw scale of MSE and runtime. Normalizing either metric before multiplication could introduce bias, as it implicitly assigns weights and may distort the true impact of each component.
Training.
In this experiment, compression and decompression are trained on a single device. For real deployment, one option is centralized training with distributed inference, meaning that the model is trained in the cloud and then the model or parameters are transmitted to edge or sensor devices for compression and decompression. Another option is distributed training with distributed inference, where models are trained on different sensors or edge devices and both compression and decompression are performed locally on those devices.
6.2. CDPI Analysis
Based on the multiply of MSE and runtime results for all models (see details in Appendix A.5 and A.6), CPDI results are presented in Table 2. The compression–decompression integrations achieve lower combined index values compared to the representative baseline models across different datasets in most cases. These results suggest that the compression–decompression framework has the potential to achieve better overall performance by reducing prediction time and lowering prediction errors with the high compression ratio.
From the perspective of different datasets, those with clear seasonal patterns, such as Electricity, tend to exhibit lower combined index values. This indicates that such datasets either require less computational time or result in lower prediction errors after processing, leading to better overall performance. Conversely, datasets with less pronounced seasonal patterns tend to have higher combined index values, reflecting increased prediction errors or longer computational times.
Table 3. Training time (in seconds) under different train horizons
Method channel=5 channel=10 channel=20 channel=40
PCDF-MLP 0.17080 0.14992 0.17723 0.58769
PCDF-Trans 0.25910 0.24673 0.27429 0.67120
PCDF-Linear 0.16091 0.17340 0.20060 0.48645
PatchTST 0.63854 0.92027 1.47746 3.64563
AutoFormer 1.53149 1.52916 1.48094 2.48258
HDMixer 0.88682 1.29958 2.06925 4.35996
Dlinear 0.07111 0.12328 0.22287 0.80155
6.3. Runtime and Accuracy Analysis
Figure 4 shows the variance of MSE. The integrations exhibit lower variance across different datasets compared to their corresponding base models, indicating that our framework provides better stability for the underlying predictors.
In terms of overall MSE and runtime values, the integrations generally achieve lower values than their base models, except for the PCDP-MLP and its corresponding base model TSMixer in the MSE metric. This suggests that our framework can effectively reduce runtime, decrease MSE, and improve overall performance compared to other methods built upon the same predictors, despite the exception observed with TSMixer.
Table 4. Compression-decompression framework ablation study via MSE (values rounded to 4 decimal places). Abbreviations: Comp. = Compression, Decomp. = Decompression, Decode. = Decoder, Encode. = Encoder.
@@ 6.3. Runtime and Accuracy Analysis
Comp.-Decomp. 0.1800 0.2153 0.1073 0.2172 0.2389
Rand.-Decomp. 0.1983 0.2595 14.8035 18.5596 0.2533
Comp.-Decode. 0.2132 1257.9678 0.2435 9.8324 0.5709
Encode.-Decomp. 0.3232 5.4924 143.5187 1.9980 12.7464
Encode.-Decode. 0.1546 1.5364 0.1737 1.6343 0.3028
6.4. Training Time Analysis
Table 3 shows the training time of the proposed framework under the same number of epochs. The PCDF-family scales well with channel count, with slower growth in training time—PCDF-Trans (×2.6), PCDF-Linear (×3.0), and PCDF-MLP (×3.4)—compared to PatchTST (×5.7), HDMixer (×4.9), and DLinear (×11.3). Even with 40 channels, all PCDF variants finish training in less than one second, showing clear speed-ups over methods without the compression–decompression framework. These results indicate that the proposed methods become more efficient when the number of input channels increases.
6.5. Ablation Analysis
In the ablation study, a classic encoder–decoder framework is used to replace the compression and decompression components in order to justify the effectiveness of the proposed framework. In addition, a random key is used in place of the seasonal key to evaluate the effectiveness of the compression component guided by seasonal information. Therefore, the ablation study includes four variants: (1) random-key compression–decompression, (2) compression–decoder, (3) encoder–decompression, and (4) encoder–decoder.
In the random-key variant, the results vary significantly across datasets, suggesting that the random key fails to uncover the underlying predictive patterns in some datasets. This highlights the importance of using the seasonal key within the compression-decompression framework.
For the compression decoder and encoder decompression variants, since compression and decompression are co-designed and interdependent, using only one of the components leads to unstable results. This indicates that optimal performance relies on the synergy between both compression and decompression modules.